tangents and normals
find the equations of all lines having slope 2 and that are tangent to the curve y=1 / (x – 3) x is not equal to 3

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

y = \frac{1}{x-3}
Slope of tangent: y’ = 2
y' = \frac{-1}{(x-3)^2}
\frac{-1}{(x-3)^2} = 2
(x-3)^2 = -2
No real root of this equation. So no tangent of slope 2 can be drawn to the curve.